OpenStudy (anonymous):

Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form

OpenStudy (istim):

Well, we could try subbing these values into y=mx+b

OpenStudy (anonymous):

y-y1=m(x-x1) x1=-2 y1=1 hen just solve for y for the point slope form :D

OpenStudy (anonymous):

Oh, okay! It makes more sense now.